Prove: lim(2n/(n ²+1))=0
1
Expert's answer
2021-10-26T15:58:04-0400

"\displaystyle
\lim_{n \to 0}(\frac{2n}{n^2+1})\\
\text{Divide each term in the limit above by the highest power of n $(n^2)$ to obtain}\\

\lim_{n \to 0}(\frac{\frac{2}{n}}{1+\frac{1}{n^2}})\\
\text{Which is 0}"